two angles are supplementary. If one angle is doubled, the result is 69% less than the other angle. Find the angles
Answer by Theo(13342)

let x = the first angle
let y = the second angle.

if the angles are supplementary, their sum is 180 degrees.

x + y = 180

if one angle is doubled, the result is 69% less than the other angle.

let x be the angle that's doubled.
let y be the angle that is 69% less.

2x = y - .69y

simplify this to get 2x = .31y.

you have 2 equations that need to be solved simultaneously.

they are:

x + y = 180
2x = .31y

rearrange the second equation to get 2x - .31y = 0

your two equations are now:

x + y = 180
2x - .31y = 0

multiply both sides of the first equation by 2 and leave the second equation as is to get:

2x + 2y = 360
2x - .31y = 0

subtract the second equation from the first to get:

2.31y = 360

solve for y to get y = 360 / 2.31 = 155.8441558 degrees.

since x + y = 180, then x = 180 - y = 24.15584416 degrees.

2x = 48.31168831 degrees.

y - .69y = 48.31168831 degrees.

looks like the problem is solved.

the angles are 155.8441558 and 24.15584416 degrees.